I did the first two but i need to know if i am correct and then i need help in the third one.
Directions for all three: Evaluate, if possible
Problem #6
sqrt121 = my answer is 11
Problem #12
sqrt (4)/(25) = my answer is (2)/(5)
Problem #3
3sqrt - (8)/(27)
This one i have no idea....

Answer by pradhyumna.kumar@gmail.com(62)

answer is simple here 3sqrt - (8)/(27) = (-2^3/3^3)^3 = -2/3